# One-Token Rollout: Guiding Supervised Fine-Tuning of LLMs with Policy Gradient

### v1

@@ 1 Introduction
As RL becomes an increasingly popular paradigm for fine-tuning LLMs, the critical role of on-policy data has garnered significant attention (Tajwar et al., 2024; Ren & Sutherland, 2024; Shenfeld et al., 2025). Tajwar et al. (2024) has shown that on-policy sampling is crucial for RL to discover optimal policies, especially when the target behavior lies in low-probability regions of the initial model. It provides a more stable and effective learning signal by ensuring that policy updates are made in regions the model can already reach, thereby preventing drastic and potentially harmful shifts in the output distribution (Ren & Sutherland, 2024; Shenfeld et al., 2025). This suggests that the on-policy nature of RL is a key factor contributing to its superior generalization and ability to preserve pre-trained knowledge.
@@ 2.1 Supervised Fine-Tuning
The standard approach for adapting pre-trained LLMs to specific downstream tasks is Supervised Fine-Tuning. Given a dataset of prompt-response pairs, where the response $X$ is a sequence of tokens $\{x_{1},x_{2},\dots,x_{T}\}$, SFT aims to maximize the conditional probability of the ground-truth sequence. This is typically achieved by minimizing the negative log-likelihood loss, autoregressively training the model $\pi_{\theta}$ to predict the next token $x_{t}$ given the preceding context $x_{1:t-1}$:
$\mathcal{L}_{\text{SFT}}(\theta)=-\frac{1}{T}\sum_{t=1}^{T}\log\pi_{\theta}(x_{t}|x_{1:t-1}).$ (1)
@@ 2.2 Policy Gradient
Policy gradient represents a class of reinforcement learning algorithms that directly optimize a parameterized policy, $\pi_{\theta}$. In this framework, the text generation process is modeled step-by-step. At each timestep $t$, the state $s_{t}$ is the sequence of previously generated tokens $x_{1:t-1}$, and the action $a_{t}$ is the next token selected by the policy from the vocabulary.
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\left(\sum_{t=1}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\right)\left(\sum_{t=1}^{T}r(s_{t},a_{t})\right)\right],$ (3)
where $\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})$ indicates the direction in the parameter space that would be used to update the policy $\pi_{\theta}$. This direction is then weighted by the sum of all rewards in the trajectory, effectively reinforcing action sequences that lead to higher overall rewards.
@@ 3 Methodology
3.1 From Policy Gradient to One-Token Rollout
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\left[\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\cdot r(s_{t},a_{t})\right].$ (4)
To implement this, we approximate the expectation $\mathbb{E}[\cdot]$ using Monte Carlo estimation. At each timestep $t$ of the original sequence, we perform a “rollout” by sampling multiple candidate actions from the current policy. This transforms the optimization problem into a practical, sample-based loss function.
@@ 3.2 Token-Level Rollout and On-Policy Reward
Stochastic Policy for Exploration. For a given state $s_{t}$, the LLMs first compute a vector of raw, unnormalized scores for every token in the vocabulary $V$. These scores are known as logits. Let $l_{a}$ denote the logit corresponding to a specific action $a$. The model’s base policy, $\pi_{\theta}$, is typically derived by applying the softmax function directly to these logits.
To encourage exploration during the rollout, we create a new sampling policy, $\pi^{\prime}_{\theta}$, by introducing a temperature parameter $\kappa$. The sampling policy is defined as:
@@
$\mathcal{A}^{\prime}_{t}=\{a^{\prime}_{t,j}\}_{j=1}^{K},\quad\text{where each }a^{\prime}_{t,j}\sim\pi^{\prime}_{\theta}(\cdot|s_{t}).$ (6)
@@
Here, $\beta$ is a hyperparameter where $\beta<1$. A reward of 1 is given for “rediscovering” the ground-truth token, while a lesser reward $\beta$ is given for all other tokens. We finally set $\beta=-0.1$ for our main experiments based on the ablation study detailed in Section 4.4.
@@ 3.3 The OTR Objective Function
$\mathcal{L}_{\text{OTR}}^{t}(\theta)=-\frac{1}{K}\sum_{j=1}^{K}\left[\text{sg}\left(R(a^{\prime}_{t,j},x_{t})\right)\cdot\log\pi_{\theta}(a^{\prime}_{t,j}|s_{t})\right],$ (8)
where $\pi_{\theta}$ is the model’s original, non-temperature-scaled policy, and $\text{sg}(\cdot)$ is the stop-gradient operator. Given our defined reward function, we can decompose this loss. Let $N_{gt}=\sum_{j=1}^{K}\mathbb{I}(a^{\prime}_{t,j}=x_{t})$ be the count of times the ground-truth token was sampled. The loss function simplifies to its final form:
$\mathcal{L}_{\text{OTR}}^{t}(\theta)=-\frac{1}{K}\left[N_{gt}\log\pi_{\theta}(x_{t}|s_{t})+\beta\sum_{j\text{ s.t. }a^{\prime}_{t,j}\neq x_{t}}\log\pi_{\theta}(a^{\prime}_{t,j}|s_{t})\right].$ (9)
@@
3.4 Comparison with Dynamic Fine-Tuning
The OTR framework can be seen as a generalization of DFT. This relationship becomes clear when we consider the special case of our OTR objective where the hyperparameter $\beta=0$. In this scenario, the second term in Equation 9, which penalizes incorrect samples, vanishes. Then the OTR loss can be formulated as:
$\mathcal{L}_{\text{OTR}}^{t}(\theta)|_{\beta=0}=-\frac{N_{gt}}{K}\log\pi_{\theta}(x_{t}|s_{t}),$ (12)
where $N_{gt}/K$ represents the empirical frequency of sampling the ground-truth token during the rollout. This frequency is a direct Monte Carlo approximation of the ground-truth token’s probability, i.e., $\frac{N_{gt}}{K}\approx\pi_{\theta}(x_{t}|s_{t})$. Thus, when $\beta=0$, the OTR objective is functionally equivalent to the DFT objective, as both methods effectively weight the loss of the ground-truth token by its estimated probability.
4 Experiments
4.1 Experiment Settings
Dataset and Models. We conduct experiments on the OpenR1-Math-220k dataset (OpenR1 Team, 2025), which consists of 220,000 mathematical problems with detailed reasoning traces. These traces are generated by the DeepSeek R1 model (DeepSeek-AI et al., 2025) for problems originating from the NuminaMath-1.5 dataset (LI et al., 2024). To efficiently manage computational resources while ensuring data quality, we randomly sample a subset of 5,000 instances for our training set. All selected instances have reasoning traces with lengths under 8192 tokens, and their lengths are approximately uniformly distributed across different intervals. We utilize a suite of powerful and contemporary open-source LLMs as base models. Specifically, we conduct our experiments on the following models: Qwen2.5-3B (Qwen Team, 2024), Qwen2.5-7B (Qwen Team, 2024), Qwen3-4B-Base (Qwen Team, 2025), and Qwen3-8B-Base (Qwen Team, 2025).
Training Details. Our implementation is built upon the Verl framework, and to ensure a fair comparison, both our proposed OTR algorithm and the SFT baseline are trained using identical settings. We employ the AdamW optimizer with a learning rate of $5\times 10^{-6}$. The learning rate follows a cosine decay schedule, which includes a warm-up ratio of 0.03 and decays to $1\times 10^{-6}$. For the training configuration, we use a batch size of 64 and a maximum sequence length of 10240 tokens. All models are trained for a total of 2 epochs.
@@ 4.1 Experiment Settings
Table 2: Out-of-domain performance on code generation and general reasoning benchmarks. For each model, the best result between SFT and OTR is in bold. The symbol indicates performance degradation compared to the base model.
@@ 4.2 Evaluation
Fine-tuned Model Evaluation. For the chat models fine-tuned with SFT and OTR, we use their respective chat templates and a 0-shot setting across all benchmarks. The decoding strategy is stochastic sampling with a temperature of 0.7 and a top-p of 0.8.
@@ 4.3 Results
Out-of-Domain Generalization. The advantages of OTR extend to out-of-domain tasks, as detailed in Table 2. On both code generation and general reasoning benchmarks, OTR consistently surpasses SFT in average performance across all models. This trend demonstrates that OTR effectively leverages its on-policy signal to achieve broader, more generalized capabilities that are not confined to its training domain.
From the perspective of knowledge preservation, OTR again proves to be a more generalizable algorithm. SFT underperforms its base model in 7 OOD instances, particularly showing vulnerability on SuperGPQA and MMLU-Pro with larger models. In contrast, OTR underperforms in 5 instances and shows consistent improvements on general reasoning for the Qwen3-4B model where SFT struggles. This demonstrates that OTR provides a more reliable fine-tuning approach that not only enhances target skills but also better maintains the model’s general intelligence, leading to superior overall generalization. A supplementary experiment is detailed in Appendix A.
@@ 4.4 Ablation Study
OTR
$\beta=\left\{\begin{tabular}[]{r}-1.00\\ -0.10\\ 0.00\\ 0.01\end{tabular}\right$. 83.10 63.05 26.05 25.75 6.88 2.91 37.66 35.06
@@
OTR
$\beta=\left\{\begin{tabular}[]{r}-1.00\\ -0.10\\ 0.00\\ 0.01\end{tabular}\right$. 92.15 77.75 40.60 35.68 12.09 13.33 53.75 46.48
@@
OTR
$\beta=\left\{\begin{tabular}[]{r}-1.00\\ -0.10\\ 0.00\\ 0.01\end{tabular}\right$. 57.90 49.70 53.80 7.28 19.07 36.26 20.87
@@
OTR
$\beta=\left\{\begin{tabular}[]{r}-1.00\\ -0.10\\ 0.00\\ 0.01\end{tabular}\right$. 74.20 61.10 67.65 9.91 28.65 56.51 31.69
@@
The Importance of Negative Samples. This study also provides insight into the importance of utilizing negative samples. As analyzed in Section 3.4, OTR with $\beta=0$ can be viewed as a Monte Carlo approximation of the DFT method. A direct comparison between the $\beta=-0.1$ and $\beta=0$ rows in our tables reveals that the former almost universally outperforms the latter. This result provides empirical evidence that incorporating an explicit penalty for negatively sampled tokens is a crucial component of OTR’s success, contributing to a more effective learning signal than what is offered by SFT-like formulations.
@@ 5 Related Work
Reinforcement Learning for Language Models. Recently, reinforcement learning has gained significant traction as a powerful paradigm for enhancing the capabilities of large language models (Hu, 2025; DeepSeek-AI et al., 2025; Wu et al., 2025a). The success of state-of-the-art models, which have leveraged RL-based algorithms like GRPO (Shao et al., 2024) to achieve substantial improvements in reasoning and cross-domain generalization, has catalyzed a surge of interest in these methods. The traditional approach to RL fine-tuning, reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), often relies on complex and computationally intensive algorithms like PPO (Schulman et al., 2017). The inherent instability and implementation complexity of PPO have motivated a recent wave of research focused on simplifying the RLHF pipeline. A prominent line of work, including methods like DPO (Rafailov et al., 2023) and GEPO (Wu et al., 2025b), elegantly reframes the preference learning objective to create a simple loss, eliminating the need for an explicit reward model. In a similar spirit, GPG (Chu et al., 2025b) simplifies the RL objective into a weighted maximum likelihood form, demonstrating that such a simplified approach can match the performance of more complex algorithms.
Improving Supervised Fine-tuning. While SFT is the most widely used paradigm for fine-tuning, its limitations, such as catastrophic forgetting and deviation from the pre-trained model’s distribution, are well-documented (Kumar et al., 2022; Huan et al., 2025). A major line of research aims to improve SFT by modifying its objective function. A prominent example is proximal SFT (Zhu et al., 2025), which introduces a proximal regularization term to the SFT loss to penalize divergence from the initial model’s policy. This approach is analogous to the KL-divergence constraint in PPO and helps stabilize training and preserve pre-trained knowledge. Another significant line of work seeks to enhance SFT by reformulating it through the lens of reinforcement learning, often by establishing a mathematical connection between their objectives. For instance, some studies reframe RLHF as a reward-weighted form of SFT (Du et al., 2025b), while others view SFT as an RL method with an implicit reward function (Wang et al., 2025; Qin & Springenberg, 2025). Concurrent to our work, DFT (Wu et al., 2025c) identifies an implicit inverse-probability weighting in the SFT gradient and addresses the resulting instability by re-weighting the loss for the ground-truth token with its own model probability. Although these works build a theoretical bridge between SFT and RL, they primarily focus on re-weighting the loss for the static, ground-truth expert data. In contrast, our work offers a distinct, data-centric solution. OTR moves beyond loss modification and instead transforms the training data itself into a dynamic, on-policy signal by actively sampling from the model’s current policy.

### v2

@@ 1 Introduction
As RL becomes an increasingly popular paradigm for fine-tuning LLMs, the critical role of on-policy data has garnered significant attention (Tajwar et al., 2024; Ren and Sutherland, 2024; Shenfeld et al., 2025). Tajwar et al. (2024) has shown that on-policy sampling is crucial for RL to discover optimal policies, especially when the target behavior lies in low-probability regions of the initial model. It provides a more stable and effective learning signal by ensuring that policy updates are made in regions the model can already reach, thereby preventing drastic and potentially harmful shifts in the output distribution (Ren and Sutherland, 2024; Shenfeld et al., 2025). This suggests that the on-policy nature of RL is a key factor contributing to its superior generalization and ability to preserve pre-trained knowledge.
@@ 2.1 Supervised Fine-Tuning
The standard approach for adapting pre-trained LLMs to specific downstream tasks is Supervised Fine-Tuning. Given a dataset of prompt-response pairs, where the prompt-response pair is a sequence of tokens $\{p,x_{1},x_{2},\dots,x_{T}\}$, SFT aims to maximize the conditional probability of the ground-truth sequence. This is achieved by minimizing the negative log-likelihood loss, autoregressively training the model $\pi_{\theta}$ to predict the next token $x_{t}$ given the prompt $p$ and the preceding context $x_{1:t-1}$:
$\mathcal{L}_{\text{SFT}}(\theta)=-\frac{1}{T}\sum_{t=1}^{T}\log\pi_{\theta}(x_{t}|p,x_{1:t-1}).$ (1)
@@ 2.2 Policy Gradient
Policy gradient represents a class of reinforcement learning algorithms that directly optimize a parameterized policy, $\pi_{\theta}$. In this framework, the text generation process is modeled step-by-step. At each timestep $t$, the state $s_{t}$ is the sequence of the prompt $p$ and previously generated tokens $x_{1:t-1}$, and the action $a_{t}$ is the next token selected by the policy from the vocabulary.
@@
$\begin{split}\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\bigg[\Big(\sum_{t=1}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\Big)\times\Big(\sum_{t=1}^{T}r(s_{t},a_{t})\Big)\bigg],\end{split}$ (3)
where $\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})$ indicates the direction in the parameter space that would be used to update the policy $\pi_{\theta}$. This direction is then weighted by the sum of all rewards in the trajectory, effectively reinforcing action sequences which can lead to higher overall rewards.
@@ 3 Methodology
3.1 From Policy Gradient to One Token Rollout
@@
$\begin{split}\nabla_{\theta}J(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\bigg[\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\times r(s_{t},a_{t})\bigg].\end{split}$ (4)
To implement this, we approximate the expectation $\mathbb{E}[\cdot]$ using Monte Carlo estimation. At each timestep $t$ of the original sequence, we perform a “rollout” by sampling candidate actions from the current policy. This transforms the optimization problem into a sample-based loss function.
@@ 3.2 Token-Level Rollout and On-Policy Reward
Stochastic Policy for Exploration. For a given state $s_{t}$, the LLMs first compute a vector of raw, unnormalized scores for every token in the vocabulary $V$. These scores are known as logits. Let $l_{a}$ denote the logit corresponding to a specific action $a$. The model’s base policy, $\pi_{\theta}$, is typically derived by applying the softmax function to these logits.
To encourage exploration, we create a new sampling policy, $\pi^{\prime}_{\theta}$, by introducing a temperature parameter $\kappa$. The sampling policy is defined as:
@@
$\mathcal{A}^{\prime}_{t}=\{a^{\prime}_{t,j}\}_{j=1}^{K},\quad a^{\prime}_{t,j}\sim\pi^{\prime}_{\theta}(\cdot|s_{t}).$ (6)
@@
Here, $\beta$ is a hyperparameter where $\beta<1$. A reward of 1 is given for “rediscovering” the ground-truth token, while a lesser reward $\beta$ is given for all other tokens. We finally set $\beta=-0.1$ based on the ablation study detailed in Section 4.4.
@@ 3.3 The OTR Objective Function
$\begin{split}\mathcal{L}_{\text{OTR}}^{t}(\theta)=-\frac{1}{K}\sum_{j=1}^{K}\bigg[R(a^{\prime}_{t,j},x_{t})\times\log\pi_{\theta}(a^{\prime}_{t,j}|s_{t})\bigg],\end{split}$ (8)
where $\pi_{\theta}$ is the model’s original, non-temperature-scaled policy. Given our defined reward function, we can decompose this loss. Let $N_{gt}=\sum_{j=1}^{K}\mathbb{I}(a^{\prime}_{t,j}=x_{t})$ be the count of times the ground-truth token was sampled. The loss function simplifies to its final form:
$\begin{split}\mathcal{L}_{\text{OTR}}^{t}(\theta)=-\frac{N_{gt}}{K}\log\pi_{\theta}(x_{t}|s_{t})-\frac{\beta}{K}\sum_{j\text{ s.t. }a^{\prime}_{t,j}\neq x_{t}}\log\pi_{\theta}(a^{\prime}_{t,j}|s_{t}).\end{split}$ (9)
@@
Olmo3-7B Base 71.20 47.00 16.00 9.60 6.70 0.00 12.50 23.29
SFT 79.95 61.00 27.38 22.33 10.41 17.51 41.25 37.12
OTR 80.73 62.10 31.60 21.75 17.91 18.11 47.81 40.00
4 Experiments
4.1 Experiment Settings
Dataset and Models. We conduct experiments on the OpenR1-Math-220k dataset (OpenR1 Team, 2025), which consists of 220,000 mathematical problems with detailed reasoning traces. These traces are generated by the DeepSeek R1 model (DeepSeek-AI et al., 2025) for problems originating from the NuminaMath-1.5 dataset (LI et al., 2024). To efficiently manage computational resources while ensuring data quality, we randomly sample a subset of 5,000 instances for our training set. All selected instances have reasoning traces with lengths under 8192 tokens, and their lengths are approximately uniformly distributed across different intervals. We utilize a suite of powerful and contemporary open-source LLMs as base models. Specifically, we conduct our experiments on the following models: Qwen2.5-3B (Qwen Team, 2024), Qwen2.5-7B (Qwen Team, 2024), Qwen3-4B-Base (Qwen Team, 2025), Qwen3-8B-Base (Qwen Team, 2025), and Olmo3-7B (Olmo, 2025). Given that OTR maintains a computational footprint highly comparable to SFT, exhibiting similar training durations and peak GPU memory usage, we select SFT as our primary baseline for comparison. We provide a comprehensive computational efficiency analysis in Appendix A.
Our implementation is built upon the Verl framework, and to ensure a fair comparison, both our OTR algorithm and the SFT baseline are trained using identical settings. We employ the AdamW optimizer with a learning rate of $5\times 10^{-6}$. The learning rate follows a cosine decay schedule, which includes a warm-up ratio of 0.03 and decays to $1\times 10^{-6}$. For the training configuration, we use a batch size of 64 and a maximum sequence length of 10240 tokens. All models are trained for 2 epochs.
Table 2: Out-of-domain performance on code generation and general reasoning benchmarks. For each model, the best result is shown in bold. The symbol indicates performance degradation compared to the base model.
@@ 4.2 Evaluation
Fine-Tuned Model Evaluation. For models fine-tuned with SFT and OTR, we use their respective chat templates and a 0-shot setting across all benchmarks. The decoding strategy is stochastic sampling with a temperature of 0.7 and a top-p of 0.8.
@@ 4.3 Results
Out-of-domain Generalization. The advantages of OTR extend to OOD tasks, as detailed in Table 2. On both code generation and general reasoning benchmarks, OTR consistently surpasses SFT in average performance across all models. This trend demonstrates OTR effectively leverages its on-policy signal to achieve more generalized capabilities that are not confined to its training domain.
From the perspective of knowledge preservation, OTR again proves to be a more generalizable algorithm. SFT underperforms its base model in 7 OOD instances, particularly showing vulnerability on SuperGPQA and MMLU-Pro with larger models. In contrast, OTR underperforms in 5 instances and shows consistent improvements on general reasoning for the Qwen3-4B model where SFT struggles. This demonstrates that OTR provides a more reliable fine-tuning approach that not only enhances target skills but also better maintains the model’s general intelligence, leading to superior overall generalization. A supplementary experiment is detailed in Appendix B.
@@ 4.4 Ablation Study
OTR $\beta=\left\{\begin{tabular}[]{r}-1.00\\ -0.10\\ 0.00\\ 0.01\end{tabular}\right$. 83.10 63.05 26.05 25.75 6.88 2.91 37.66 35.06
@@
OTR $\beta=\left\{\begin{tabular}[]{r}-1.00\\ -0.10\\ 0.00\\ 0.01\end{tabular}\right$. 92.15 77.75 40.60 35.68 12.09 13.33 53.75 46.48
@@
OTR $\beta=\left\{\begin{tabular}[]{r}-1.00\\ -0.10\\ 0.00\\ 0.01\end{tabular}\right$. 57.90 49.70 53.80 7.28 19.07 36.26 20.87
@@
OTR $\beta=\left\{\begin{tabular}[]{r}-1.00\\ -0.10\\ 0.00\\ 0.01\end{tabular}\right$. 74.20 61.10 67.65 9.91 28.65 56.51 31.69
@@
The Importance of Negative Samples. This study also provides insight into the importance of utilizing negative samples. As analyzed in Appendix C, OTR with $\beta=0$ can be viewed as a Monte Carlo approximation of the DFT method. A direct comparison between the $\beta=-0.1$ and $\beta=0$ rows in our tables reveals that the former almost universally outperforms the latter. This result provides empirical evidence that incorporating an explicit penalty for negatively sampled tokens is a crucial component of OTR’s success, contributing to a more effective learning signal than what is offered by SFT-like formulations.
@@ 5 Related Work
RL for LLMs. Recently, RL has gained significant traction as a powerful paradigm for enhancing the capabilities of large language models (Hu, 2025; DeepSeek-AI et al., 2025; Wu et al., 2025a). The success of state-of-the-art models, which have leveraged RL-based algorithms like GRPO (Shao et al., 2024) to achieve substantial improvements in reasoning and cross-domain generalization, has catalyzed a surge of interest in these methods. The traditional approach to RL fine-tuning, reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), often relies on complex and computationally intensive algorithms like PPO (Schulman et al., 2017). The inherent instability and implementation complexity of PPO have motivated a recent wave of research focused on simplifying the RLHF pipeline. A prominent line of work, including methods like DPO (Rafailov et al., 2023) and GEPO (Wu et al., 2025b), elegantly reframes the preference learning objective to create a simple loss, eliminating the need for an explicit reward model. In a similar spirit, GPG (Chu et al., 2025b) simplifies the RL objective into a weighted maximum likelihood form, demonstrating that such a simplified approach can match the performance of more complex algorithms.
Improving SFT. While SFT is the most widely used paradigm for fine-tuning, its limitations, such as catastrophic forgetting and deviation from the pre-trained model’s distribution, are well-documented (Kumar et al., 2022; Huan et al., 2025). A major line of research aims to improve SFT by modifying its objective function. A prominent example is proximal SFT (Zhu et al., 2025), which introduces a proximal regularization term to the SFT loss to penalize divergence from the initial model’s policy. This approach is analogous to the KL-divergence constraint in PPO and helps stabilize training and preserve pre-trained knowledge. Another significant line of work seeks to enhance SFT by reformulating it through the lens of reinforcement learning, often by establishing a mathematical connection between their objectives. For instance, some studies reframe RLHF as a reward-weighted form of SFT (Du et al., 2025b), while others view SFT as an RL method with an implicit reward function (Wang et al., 2025; Qin and Springenberg, 2025). Concurrent to our work, DFT (Wu et al., 2025c) identifies an implicit inverse-probability weighting in the SFT gradient and addresses the resulting instability by re-weighting the loss for the ground-truth token with its own model probability. Although these works build a theoretical bridge between SFT and RL, they primarily focus on re-weighting the loss for the static, ground-truth expert data. In contrast, our work offers a distinct, data-centric solution. OTR moves beyond loss modification and instead transforms the training data itself into a dynamic, on-policy signal by actively sampling from the model’s current policy.